# Synergy: Towards On-Body AI via Tiny AI Accelerator Collaboration on Wearables

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
The advent of tiny artificial intelligence (AI) accelerators, such as the Analog MAX78000 (Max78000, [n. d.]), MAX78002 (Max78002, [n. d.]), Google Coral Micro (Coral Micro, [n. d.]) and GreenWaves GAP-9 (GAP8/GAP9, [n. d.]) has brought AI closer to us than ever before, offering reduced latency, low power cost, and improved privacy. These accelerators, designed for microcontrollers (MCUs) with small form factors (e.g., MAX78000: 8mm×$\times$8mm in Figure 1), are becoming integrated into wearable devices recently (Balaji and Peh, 2023; OmniBuds, [n. d.]; Shift Moonwalkers, [n. d.]; Hearables with GAP9, [n. d.]), e.g., smart earbuds, patches, watches, glasses, wristband, and shoes. This integration transforms wearable devices from smartphone-dependent accessories—used merely for data collection or notification alerts—into AI-capable devices.
With the proliferation of wearable devices, it is natural to expect that personal computing environments on the body will form a network of these AI-capable wearables. This enables a new class of on-body AI apps to emerge, which enhance app functionality by leveraging on-device AI capabilities with diverse sensing capabilities and interaction methods available across wearables. These apps, running on wearable devices, continuously monitor a variety of user contexts and proactively provide context-aware services to users directly via diverse wearable interfaces. Figure 1 shows an example scenario with three concurrent on-body AI apps: (i) memory augmentation detects greeting words using a smartwatch microphone and captures nearby faces using glasses-mounted cameras, (ii) attention alert monitors surrounding visual events through glasses and provides haptic alerts on a ring, and (iii) personal fitness coach analyzes exercise routines and vital signs on a smartwatch and smart shoes, providing auditory feedback via earbuds.
Figure 1. Wearable computing powered by tiny AI accelerators, simultaneously running on-body AI apps.
@@ 1. Introduction
However, it is not straightforward to realize these benefits. In the current paradigm for multi-wearable programming, task-device assignment decisions are often made at the development time. These individual app-level decisions may not be optimal at runtime due to the dynamic nature of wearable devices and dependencies on resource usage between concurrent apps. One may consider existing model partitioning techniques (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021). While they support dynamic model splitting, they mainly focus on optimizing the splitting decision for a single AI model, thereby lacking a holistic view that considers (i) interdependency among different tasks—sensing, model inference, interaction on distributed devices—within an app and (ii) resource conflict across concurrent apps. Additionally, no existing partitioning work has been built for tiny AI accelerators yet, making it difficult to directly adopt them for on-body AI apps.
@@
We prototyped Synergy on two tiny AI accelerator platforms, MAX78000 and MAX78002. We compare Synergy with 7 baselines including state-of-the-art model partitioning techniques (Kang et al., 2017; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Jeong et al., 2018; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021). Our extensive evaluation with 8 AI models shows that Synergy consistently outperforms the baselines, with on average 23.0$\times$ throughput gain across various scenarios, while reducing latency by 73.9% and power consumption by 15.8%. Our in-depth experiments also show that Synergy effectively adapts to various runtime environment changes: the number of devices, the number of pipelines, heterogeneous device resources, different source and target mapping, and different objectives.
2. Background & Motivation
2.1. Tiny AI Accelerators
The integration of AI accelerators into MCUs represents a significant move towards distributed, on-device AI, ensuring enhanced user privacy and minimal latency. Although a number of tiny-scale accelerators have been proposed recently, very few products are commercially available and provide access and control over their underlying operations. In this paper, we chose Analog MAX78000 (Max78000, [n. d.]) and MAX78002 (Max78002, [n. d.]) as our primary platforms since they are the most widely used tiny AI accelerator research platforms (Moosmann et al., 2023; Rüegg et al., 2023; Bakar et al., 2022; Caronti et al., 2023; Moss et al., 2022) owing to the disclosed hardware details and open-source tools, enabling in-depth analysis and modification of operations.
@@ 2.1. Tiny AI Accelerators
Figure 2. Performance comparison between AI accelerator (MAX78000) and MCUs (MAX32650 and STM32F7).
Recent benchmark study (MAX78000 Performance, [n. d.]; Moss et al., 2022) quantifies the MAX78000’s superior performance in terms of latency and energy cost. Figure 2 shows MAX78000 significantly outperforms a conventional MCU, MAX32650 with Cortex-M4 at 120 MHz (MAX32650, [n. d.]) and even a high-performance MCU, STM32F7 with Cortex-M7 at 216 MHz (STM32F7, [n. d.]) in key AI models. Latency for keyword spotting (KWS) is reduced to 2.0 ms compared to 350 ms and 123 ms for MAX32650 and STM32F7, respectively. Energy efficiency is similarly enhanced, with MAX78000 consuming merely 0.40 mJ for face detection (FaceID), in contrast to 42.1 mJ and 464 mJ consumed by MAX32650 and STM32F7.
2.2. AI Accelerator Collaboration
@@
Figure 3. Comparison between (a) AI accelerator collaboration via Synergy and (b) phone offloading.
Figure 4. Comparison of Synergy and phone offloading.
2.2.1. Why not offloading?
@@
Offloading AI model execution to a smartphone might seem advantageous due to its higher processing capability. However, this approach can be inefficient for on-body AI apps because (i) they require continuous sensing and inference for situational services and (ii) services are directly provided through wearable devices. As shown in Figure 3, offloading incurs additional data communication, resulting in increased latency and energy costs. These costs are unnecessary when apps run solely on wearables. Continuous transmission of raw sensor data to a smartphone can also cause communication and energy bottlenecks, affecting both smartphones and wearables. We compared the total throughput of model execution using Workload 1 and 2 from Section 6.2. By eliminating unnecessary links, Synergy improves throughput by 57.7$\times$ and 28.8$\times$ compared to smartphone offloading, as shown in Figure 4. Despite frequent model execution, Synergy consumes less or comparable power due to avoiding energy-intensive data communication.
3. Synergy Design
3.1. Holistic Orchestration for AI Accelerator Collaboration
3.1.1. Limitations of Existing Methods
In the current paradigm for multi-wearable programming, task-device assignments are predetermined in the app logic at development time. This app-level assignment decision has limitations in environments with concurrent AI apps on heterogeneous wearables with varying availability. For example, a memory augmentation app using a smartwatch’s AI accelerator would underperform or fail if other apps compete for the same accelerator or if the smartwatch’s battery depletes, even when other devices have available accelerators. Leveraging multiple AI accelerators is also challenging because developers lack visibility into available accelerators and their capabilities at development time.
Figure 5. Comparison of decision approaches.
Existing model partitioning techniques (Kang et al., 2017; Jeong et al., 2018; Hu et al., 2019; Zhang et al., 2020; Laskaridis et al., 2020; Wang et al., 2021; Banitalebi-Dehkordi et al., 2021) designed to split AI models for offloading to powerful devices (smartphones or cloud) could be adapted for distributed on-body AI accelerators. However, these techniques primarily focus on optimizing the splitting decision for a single model. Applying these techniques to on-body AI apps presents challenges because they lack a holistic view considering (i) resource dependency among concurrent apps and (ii) interdependency between model inference and other tasks such as sensing and interaction. Additionally, no existing work is designed for partitioning models onto tiny AI accelerators, thus lacking consideration of their unique characteristics for performance modeling.
@@ 3.1.1. Limitations of Existing Methods
Independent decision: Existing partitioning methods split each concurrent model individually and deploy the chunks to assigned devices. This independent approach can cause resource conflicts among models, leading to performance degradation and OOR failures, especially with memory-constrained AI accelerators. For instance, as illustrated in Figure 5 (a), if two apps deploy their model chunks to the same device ($d_{1}$)—the full model of $p_{1}$’s $m_{1}$ and the partial model of $p_{2}$’s $m_{2}$—it can cause OOR on $d_{1}$ due to memory constraints.
Model-centric joint decision: Extending these approaches to jointly consider the memory capacity of split models and filter out unsupported cases can mitigate resource conflicts. However, this model-centric view still results in suboptimal performance. An app’s end-to-end pipeline includes sensing and interaction tasks before and after model inference. Data communication costs for transmitting data from the sensing device (source) to the model execution device, and then to the destination device (target), impact the app’s latency and throughput. Figure 5 (b) illustrates this. Although the split decision for $m_{1}$ and $m_{2}$ on $p_{1}$ and $p_{2}$ avoids resource conflicts, it can still cause additional communication overhead when execution is far from the source and target devices.
3.1.2. Our Approach: AI Accelerator Collaboration with a Holistic View
We propose shifting from app-level independent decision to system-driven collaboration for environments with concurrent AI apps on distributed AI accelerator-equipped wearables. The key is to decouple task-device assignment from app logic, giving the runtime system visibility and fine-grained control over AI accelerator collaboration. This allows runtime optimization of collaboration planning, leveraging direct knowledge of concurrent workloads, available resources, and dependencies. Synergy offers developers a device-agnostic programming interface (§4.2), enabling them to define the end-to-end pipeline without specifying target devices. At runtime, Synergy dynamically assigns tasks to devices to ensure best-effort performance.
The core of runtime orchestration is to exploit collaboration opportunities by (i) flexibly assigning tasks to available devices based on task requirements and device capabilities, and (ii) splitting AI models across distributed AI accelerators. We adopt a holistic view to address resource dependency between tasks within an app and between concurrent apps. For each collaboration opportunity, we abstract resource use by all concurrent apps and their tasks over distributed devices into a holistic collaboration plan. We then identify the best collaboration case based on the system objective. Figure 5 (c) shows an example of holistic collaboration by Synergy. Synergy reduces end-to-end latency by placing source and target devices closer to model execution devices on top of the best model split decision, thereby enhancing throughput.
3.2. System Scope
@@
Target metric: To execute multiple models across distributed devices, various system-wide objectives can be considered, such as maximizing inference throughput, minimizing latency, and reducing energy consumption. This paper focuses on maximizing overall system-wide model inference throughput, a key quality of service metric in AI apps. To ensure fairness among models, we merge multiple app pipelines into a unified one and maximize its execution per second. Simply maximizing total throughput—the number of model executions per second—could let lower-latency models monopolize resources, causing fairness issues. This target metric can be replaced with other objectives, such as minimizing latency or energy costs. Results are reported in §6.3.4.
4. Synergy Runtime
4.1. Overview
Figure 6 illustrates the operational flow of Synergy.
Figure 6. An illustration of Synergy’s operational flow.
App development & installation stage: On-body AI apps are developed using the device-agnostic programming interfaces provided by Synergy (§4.2). Upon app installation, its end-to-end pipeline is maintained on a moderator, and tasks are dynamically distributed onto wearable devices based on orchestration decisions.
Orchestration stage: When running apps or available devices change, Synergy triggers holistic orchestration. It generates holistic collaboration plans, mapping all tasks in concurrent apps to devices (§4.3), by matching task requirements with device capabilities and splitting model tasks across distributed AI accelerators. Synergy then evaluates the potential throughput of each plan and selects the one with the maximum throughput. Two important aspects are:
  • A complete search for the optimal plan introduces intractable computation complexity, especially in resource-constrained MCUs. To reduce the search space effectively, we propose a progressive search space reduction method (§4.4), scaling down computational complexity from exponential to linear with the number of pipelines, while keeping the selected plan close to optimal.
  • Evaluating the potential throughput of candidates is crucial for plan selection. We devise a novel latency estimation model to estimate AI model execution time on tiny AI accelerators. Based on this, we build an online throughput estimation method (§4.5) to estimate the latency and throughput of a collaboration plan.
Post-deployment stage: As we combine multiple pipelines into a holistic collaboration plan, the execution cycle completes when all pipelines finish. This can lead to idle times for tasks that can be executed on parallel computation units (e.g., MCU, AI accelerator, wireless chip). To improve throughput at runtime, we propose adaptive task parallelization (§4.6), maximizing parallel execution opportunities across distributed computation units and reducing idle times.
4.2. Device-Agnostic Programming Interface

### v3

@@ Abstract
Index Terms:
On-body AI, tiny AI accelerator, Accelerator composition, wearables.
I Introduction
The advent of tiny artificial intelligence (AI) accelerators, such as the Analog MAX78000 [1], MAX78002 [2], Google Coral Micro [3] and GreenWaves GAP-9 [4] has brought AI closer to us than ever before, offering reduced latency, low power cost, and improved privacy. These accelerators, designed for microcontrollers (MCUs) with small form factors (e.g., MAX78000: 8mm×$\times$8mm in Figure 1), are becoming integrated into wearable devices recently [5, 6, 7, 8], e.g., smart earbuds, patches, watches, glasses, wristband, and shoes. This integration transforms wearable devices from smartphone-dependent accessories—used merely for data collection or notification alerts—into AI-capable devices.
With the proliferation of wearable devices, it is natural to expect that personal computing environments on the body will form a network of these AI-capable wearables [9]. This enables a new class of on-body AI apps to emerge, which enhance app functionality by leveraging on-device AI capabilities with diverse sensing capabilities and interaction methods available across wearables. These apps, running on wearable devices, continuously monitor a variety of user contexts and proactively provide context-aware services to users directly via diverse wearable interfaces. Figure 1 shows an example scenario with three concurrent on-body AI apps: (i) memory augmentation detects greeting words using a smartwatch microphone and captures nearby faces using glasses-mounted cameras, (ii) attention alert monitors surrounding visual events through glasses and provides haptic alerts on a ring, and (iii) personal fitness coach analyzes exercise routines and vital signs on a smartwatch and smart shoes, providing auditory feedback via earbuds.
Figure 1: Wearable computing powered by tiny AI accelerators, simultaneously running on-body AI apps.
@@ I Introduction
However, it is not straightforward to realize these benefits. In the current paradigm for multi-wearable programming, task-device assignment decisions are often made at the development time. These individual app-level decisions may not be optimal at runtime due to the dynamic nature of wearable devices and dependencies on resource usage between concurrent apps. One may consider existing model partitioning techniques [10, 11, 12, 13, 14, 15, 16]. While they support dynamic model splitting, they mainly focus on optimizing the splitting decision for a single AI model, thereby lacking a holistic view that considers (i) interdependency among different tasks—sensing, model inference, interaction on distributed devices—within an app and (ii) resource conflict across concurrent apps. Additionally, no existing partitioning work has been built for tiny AI accelerators yet, making it difficult to directly adopt them for on-body AI apps. Note that such dynamic challenges can be addressed by sending raw sensor and offloading all processing tasks to smartphones, which can be assumed to be always available. However, this incurs high communication costs to both wearable devices and the smartphone, thereby leading to increased power consumption and degraded inference throughput of AI models, as discussed in §II-B.
@@
We prototyped Synergy on two tiny AI accelerator platforms, MAX78000 and MAX78002. We compare Synergy with 7 baselines including state-of-the-art model partitioning techniques [10, 11, 12, 13, 14, 15, 16]. Our extensive evaluation with 8 AI models shows that Synergy consistently outperforms the baselines, with on average 23.0$\times$ throughput gain across various scenarios, while reducing latency by 73.9% and power consumption by 15.8%. Our in-depth experiments also show that Synergy effectively adapts to various runtime environment changes: the number of devices, the number of pipelines, heterogeneous device resources, different source and target mapping, and different objectives.
II Background & Motivation
II-A Tiny AI Accelerators
The integration of AI accelerators into MCUs represents a significant move towards distributed, on-device AI, ensuring enhanced user privacy and minimal latency. Although a number of tiny-scale accelerators have been proposed recently, very few products are commercially available and provide access and control over their underlying operations. In this paper, we chose Analog MAX78000 [1] and MAX78002 [2] as our primary platforms since they are the most widely used tiny AI accelerator research platforms [17, 18, 19, 20, 21, 22, 23, 24, 25] owing to the disclosed hardware details and open-source tools, enabling in-depth analysis and modification of operations.
@@ II-A Tiny AI Accelerators
Figure 2: Performance comparison between AI accelerator (MAX78000) and MCUs (MAX32650 and STM32F7).
Recent benchmark study [26, 21] quantifies the MAX78000’s superior performance in terms of latency and energy cost. Figure 2 shows MAX78000 significantly outperforms a conventional MCU, MAX32650 with Cortex-M4 at 120 MHz [27] and even a high-performance MCU, STM32F7 with Cortex-M7 at 216 MHz [28] in key AI models. Latency for keyword spotting (KWS) is reduced to 2.0 ms compared to 350 ms and 123 ms for MAX32650 and STM32F7, respectively. Energy efficiency is similarly enhanced, with MAX78000 consuming merely 0.40 mJ for face detection (FaceID), in contrast to 42.1 mJ and 464 mJ consumed by MAX32650 and STM32F7.
II-B AI Accelerator Collaboration
@@
Figure 3: Comparison between (a) AI accelerator collaboration via Synergy and (b) phone offloading.
Figure 4: Comparison of Synergy and phone offloading.
II-B1 Why not offloading?
Offloading AI model execution to a smartphone might seem advantageous due to its higher processing capability. In particular, this would be true for conventional mobile apps where model execution occurs intermittently, mostly upon a user’s request, and the service with the final inference output is provided on a smartphone. However, this approach can be inefficient for on-body AI apps because (i) they require continuous sensing and inference for situational services and (ii) services are directly provided through wearable devices. Figure 3 shows an illustrative comparison when the three apps introduced in §I are concurrently running. While the AI accelerator collaboration can support the concurrent execution of three apps with four device-to-device (d2d) communications as shown in Figure 3 (a), smartphone offloading requires seven d2d communications as shown in Figure 3 (b), as the data must pass through a smartphone regardless of the data path. Thus, such offloading incurs additional data communication and results in increased latency and energy costs. These additional transmission costs are unnecessary when apps run solely on wearables.
@@ II-B1 Why not offloading?
Continuous transmission of raw sensor data to a smartphone also causes communication and energy bottlenecks, affecting both smartphones and wearables. We compared the total throughput of model execution using Workload 1 and 2 from Section VI-B. By eliminating unnecessary links, Synergy improves throughput by 57.7$\times$ and 28.8$\times$ compared to smartphone offloading, as shown in Figure 4. Despite frequent model execution, Synergy consumes less or comparable power due to avoiding energy-intensive data communication.
Based on these findings, we can conclude that offloading (sending raw sensor data to a smartphone) is not effective when (i) the size of raw sensor data is larger than that of the intermediate or final results, (ii) the final results need to be delivered to wearable devices for the service, and (iii) the model needs to be executed frequently for situational and proactive services, which is the case for on-body AI apps.
III Synergy Design
Figure 5: Comparison of decision approaches.
III-A Limitations of Existing Approaches
In the current multi-wearable programming paradigm, task-device assignments are predetermined at development time, limiting adaptability in dynamic environments with concurrent AI apps and heterogeneous wearables. This app-level assignment decision poses several challenges:
  • Lack of adaptability to resource contention: If a memory augmentation app relies on a smartwatch’s AI accelerator, it may underperform or fail due to competition from other apps or battery depletion, even when other devices have available accelerators.
  • Limited visibility into available resources: Developers lack real-time awareness of available AI accelerators and their capabilities, making it difficult to leverage multiple devices efficiently.
Existing model partitioning techniques [10, 14, 11, 12, 13, 15, 16] primarily focus on offloading computation to powerful devices (e.g., smartphones or cloud). While they could be adapted for distributed on-body AI accelerators, they have fundamental limitations:
  • Single-model optimization: These techniques optimize splitting decisions for individual models, without considering: (i) Resource dependencies between concurrent apps. (ii) Interdependencies between model inference and other tasks (e.g., sensing and interaction).
  • Incompatibility with tiny AI accelerators: No existing method is specifically designed to partition models onto tiny AI accelerators, leading to inaccurate performance modeling in resource-constrained environments.
@@ III-A Limitations of Existing Approaches
  • Independent decision-making: Existing partitioning methods split and deploy each model independently, ignoring conflicts among models. This can lead to resource contention and out-of-resource (OOR) failures, particularly for memory-constrained accelerators. Example: As shown in Figure 5 (a), if two apps deploy model chunks to the same device ($d_{1}$), the full model of $p_{1}$’s $m_{1}$ and the partial model of $p_{2}$’s $m_{2}$ may exceed $d_{1}$’s memory capacity, leading to OOR failures.
  • Model-centric joint decision: Some methods extend model partitioning to account for memory constraints, filtering out unsupported cases. However, this model-centric view still remains suboptimal, as it ignores sensing and interaction costs in an app’s end-to-end pipeline. Example: As illustrated in Figure 5 (b), even when split models of $m_{1}$ and $m_{2}$ avoid memory conflicts, long-distance execution between source and target devices increases communication overhead, degrading overall performance.
These limitations highlight the need for a holistic, system-driven approach that considers concurrent workloads, device heterogeneity, and end-to-end execution costs.
III-B Our Approach: AI Accelerator Collaboration with a Holistic View
We propose shifting from app-level independent decision-making to system-driven collaboration in environments with concurrent AI applications on distributed AI accelerator-equipped wearables. Our approach introduces the following key innovations:
  • Decoupled task-device assignment: Existing approaches tightly couple task execution with app logic, limiting flexibility. We decouple task-device assignment, allowing the runtime system to gain system-wide visibility and fine-grained control over AI accelerator collaboration.
  • Device-agnostic programming interface: Synergy provides a programming model (§IV-B) that enables developers to define end-to-end AI pipelines without specifying target devices. This abstraction facilitates dynamic task placement, adapting execution based on real-time resource availability.
  • Holistic collaboration and runtime optimization: Unlike existing methods that optimize resource allocation per app, Synergy takes a holistic approach, considering both intra-app and inter-app dependencies. It constructs a system-wide collaboration plan by leveraging direct knowledge of concurrent workloads, resource availability, and task dependencies. The runtime dynamically assigns tasks to devices based on their capabilities and splits AI models across distributed AI accelerators to maximize efficiency.
  • End-to-end performance optimization: Beyond model partitioning, Synergy optimizes task placement by minimizing communication overhead between source, execution, and target devices. This improves throughput and reduces overall latency. Figure 5 (c) illustrates an example of this holistic collaboration.
By adopting a system-driven approach, Synergy overcomes the limitations of independent per-app optimization, improving efficiency, reducing latency, and enhancing AI accelerator utilization in multi-app distributed AI environments.
III-C System Scope
@@
Target metric: To execute multiple models across distributed devices, various system-wide objectives can be considered, such as maximizing inference throughput, minimizing latency, and reducing energy consumption. This paper focuses on maximizing overall system-wide model inference throughput, a key quality of service metric in AI apps. To ensure fairness among models, we merge multiple app pipelines into a unified one and maximize its execution per second. Simply maximizing total throughput—the number of model executions per second—could let lower-latency models monopolize resources, causing fairness issues. This target metric can be replaced with other objectives, such as minimizing latency or energy costs. Results are reported in §VI-C4.
Potential synergy with smartphones: While Synergy primarily focuses on tiny AI accelerators on wearable devices and their collaboration, and we show its benefits over phone offloading for on-body AI apps in §II-B, incorporating smartphones could enhance system performance and flexibility in other scenarios. In particular, a hybrid approach leveraging both wearables and smartphones can be advantageous for cases where large AI models need to run with body signals, e.g., foundation models with IMU data [29] and optical physiological signals [30]. Another example would be when the final inference outcome is served using a mobile UI on a smartphone, thereby requiring no transmission overhead from a smartphone to wearable devices. By integrating the synergy with smartphones, Synergy could extend its applicability to a broader range of AI-driven wearable apps while maintaining seamless user experiences and efficient on-body computing. We leave this as future work.
IV Synergy Runtime Technical Components
IV-A Overview
This section provides a high-level overview of our approach. We first summarize the key technical novelties that distinguish Synergy from existing solutions, followed by a step-by-step description of Synergy’s operational flow in multi-wearable environments.
IV-A1 Summary of Technical Novelties
Synergy introduces several innovations that improve performance and flexibility for on-body AI applications:
  • Device-agnostic programming interface: An abstracted interface for building on-body AI pipelines without specifying target devices at design time.
  • Holistic collaboration planning: A global approach that considers all concurrent apps together, supporting multi-model splitting and effective resource sharing.
  • Progressive search space reduction: A method to mitigate the combinatorial explosion of multi-pipeline orchestration, reducing complexity from exponential to near-linear in the number of pipelines.
  • Clock-cycle-based latency modeling: A precise inference-latency estimator that captures per-layer execution time on tiny AI accelerators by analyzing hardware clock cycles.
  • Online throughput estimation: A lightweight scheme for dynamically evaluating end-to-end performance in real time, enabling adaptive selection of the best collaboration plan.
  • Adaptive task parallelization: Mechanisms that exploit idle computation units by overlapping tasks across multiple pipelines (inter-pipeline) and consecutive runs (inter-run).
These novelties overcome limitations of per-app approaches by jointly managing on-body resources. Next, we describe how Synergy integrates these techniques into its operational flow.
IV-A2 Operational Flow
Figure 6 depicts the runtime stages of Synergy:
Figure 6: An illustration of Synergy’s operational flow.
App development & installation stage: On-body AI apps are developed using the device-agnostic programming interfaceIV-B). Developers only need to specify the pipeline’s logical tasks without worrying about which wearables will execute them. Once the app is installed, the pipeline description is registered with a moderator that manages device assignments.
Orchestration stage: When running apps or available devices change, Synergy triggers holistic orchestration by generating holistic collaboration plans (§IV-C). It maps all tasks in concurrent apps to devices by matching task requirements with device capabilities and potentially splitting model tasks across distributed AI accelerators. Because exhaustively searching for the best split and device assignment can be computationally intractable on resource-constrained MCUs, Synergy applies progressive search space reductionIV-D) to prune infeasible or suboptimal plans. Once feasible candidates are formed, Synergy employs a novel latency estimation model and an online throughput estimation method (§IV-E) to select the plan expected to yield the highest overall throughput.
Post-deployment stage: As multiple pipelines execute under a single holistic collaboration plan, the entire execution cycle completes when all pipelines finish. This can create idle periods on certain computation units (e.g., MCUs, AI accelerators, wireless chips). To further boost throughput at runtime, Synergy incorporates adaptive task parallelizationIV-F), which overlaps tasks across pipelines (inter-pipeline) and across multiple runs (inter-run) to reduce idle time and improve overall system performance.
By coordinating these stages under a single, unified framework, Synergy effectively balances task distribution, model-splitting decisions, and parallelization strategies to achieve high performance and resource utilization across multiple on-body AI accelerators.
IV-B Device-Agnostic Programming Interface